OpenStudy (anonymous):

Evaluate 8 − 12 ÷ 2 + 3 1 −1 4 5

OpenStudy (anonymous):

its 1

OpenStudy (anonymous):

its 5

OpenStudy (anonymous):

8-12=-4 -4/2=-2 -2+3=1

OpenStudy (anonymous):

no its 1. it would be 5 if it was 12-8

OpenStudy (dobby1):

its -1

OpenStudy (anonymous):

seriously do that math its 1.

OpenStudy (ivettef365):

First you need to divide -12 / 2 = -6,then you have 8 -6 +3 = 5

OpenStudy (anonymous):

see it is 5

OpenStudy (anonymous):

no, you have to do the subtraction first, the rule only applies to multiplication not divsion

OpenStudy (dobby1):

\[8-12/2+3\]

OpenStudy (ivettef365):

P arenthesis E xponents M ultiplication D ivivion A dd S subtract

OpenStudy (ivettef365):

that is the order

OpenStudy (dobby1):

you divide then add then subtract google pentas

OpenStudy (anonymous):

im using a website that shows steps and it says its 5

OpenStudy (dobby1):

oops pemdas

OpenStudy (dobby1):

parenthases, exponent, multiplication, division, addition, subtraction

OpenStudy (dobby1):

oh yeah it is 5

OpenStudy (dobby1):

i made a mistake

OpenStudy (ivettef365):

that is ok, we learn from our mistakes

OpenStudy (dobby1):

yeah thanks :)